OpenStudy (anonymous):

simplify the given expression algebraically: sin(x- pi/2). is it -cosx ?

OpenStudy (anonymous):

sin(x- pi/2) = sin(-(pi/2 - x)) = - sin(pi/2 - x) = - cos(x) you are right